# RSCaMa: Remote Sensing Image Change Captioning with State Space Model

### v1

@@ III-C Comparison to State-of-the-Art
III-D Ablation Studies
III-D1 CaMa Layers
The CaMa layers consist of SD-SSM and TT-SSM. Table II validates the effectiveness of these two modules. The structure of our baseline model is similar to SD-SSM, but the key difference lies in using self-features rather than differential features to multiply with the output of bidirectional SSM, following the original design of Mamba. Experimental results indicate that SD-SSM and TT-SSM significantly contribute to improving change captioning performance. This is because SD-SSM effectively enhances the model’s understanding of spatial changes by multiplying differential features with the output of bidirectional SSM, while TT-SSM promotes effective temporal interaction of bi-temporal features through a token-wise cross-scanning manner.
@@ III-D2 Language Decoder
RSICC require transforming visual change into language. The language decoders adopted by previous methods have evolved with the development of language models, from recurrent neural networks to Transformers. As a novel language model, the performance of Mamba in cross-modal understanding is yet to be explored. Table III compares three different language decoding methods: the Mamba language model, the Transformer decoder with cross-attention, and the GPT-style decoder with causal attention only. In our experiments, for the first two methods, we prefix the visual embeddings to the language decoders to generate words one by one. For the last method, visual embeddings are integrated into the decoding process as keys (K) and values (V) of the cross-attention mechanism. The experimental results in Table III demonstrate that the Transformer decoder with cross-attention mechanism performs best in transforming visual to textual information, surpassing the simple visual embedding prefix strategy (i.e., Mamba and GPT-style decoder), thus proving the superiority of cross-attention in cross-modal understanding. In the future, it is promising to explore the combination of Mamba with novel attention mechanisms rather than the simple visual prefix strategy to enhance its cross-modal understanding ability when generating each word.
@@ III-E Qualitative Results
Fig. 4 showcases some captioning results. Each image pair is accompanied by one of the five ground-truth sentences, along with the sentences generated by the baseline model and our RSCaMa. Notably, our RSCaMa demonstrates a remarkable ability to accurately distinguish changes in the scenes, as illustrated by the first image pair. Furthermore, our RSCaMa excels in identifying the categories of changing objects, their locations, and the relationships between objects. For instance, in the second image pair, our RSCaMa accurately identifies the number of the changed house, while in the third image pair, it correctly identifies the changed building’s location in the lower right corner.
@@ IV Conclusion
In this paper, we introduce Mamba (a state space model) into the RSICC task and propose a novel approach called RSCaMa. RSCaMa iteratively processes bi-temporal features from spatial and temporal domains using multiple CaMa layers composed of SD-SSM and TT-SSM. SD-SSM uses differential features to enhance change perception, while TT-SSM promotes bitemporal interactions in a token-wise cross-scanning manner. Experimental results demonstrate the superiority of RSCaMa over existing methods, highlighting the potential of Mamba in RSICC tasks. Besides, our systematic evaluation of three language decoding schemes provides valuable insights for future research.

### v3

@@ III-D2 Language Decoder
RSICC require transforming visual change into language. The language decoders adopted by previous methods have evolved with the development of language models, from recurrent neural networks to Transformers. As a novel language model, the performance of Mamba in cross-modal understanding is yet to be explored. In Table IV, we compare three language decoders: the Mamba language model, the Transformer decoder with cross-attention, and the GPT-style decoder with causal attention only. In our experiments, for the first two methods, we prefix the visual embeddings to the language decoders to generate words one by one. For the Transformer decoder with cross-attention, visual embeddings are integrated into the decoding process as keys (K) and values (V) of the cross-attention mechanism. The experimental results in Table IV demonstrate that the Transformer decoder with cross-attention mechanism performs best in transforming visual to textual information, surpassing the simple visual embedding prefix strategy (i.e., Mamba and GPT-style decoder), thus proving the superiority of cross-attention in cross-modal understanding. In the future, it is promising to explore combining Mamba with novel attention mechanisms rather than the simple visual prefix strategy to enhance its cross-modal understanding ability when generating each word.
@@
Figure 5: Visualization of features before and after CaMa layer processing.
@@ III-E Qualitative Results
Fig. 4 showcases some captioning results. We can observe that our RSCaMa demonstrates a remarkable ability to accurately distinguish changes in the scenes, as illustrated by the first image pair. Furthermore, our RSCaMa excels in identifying the categories of changing objects, their locations, and the relationships between objects. For instance, in the second image pair, our RSCaMa accurately identifies the number of the changed houses, while in the third image pair, it correctly recognizes that the location of changed buildings is in the lower right corner.
In Fig. 5, we visualise bi-temporal features before and after CaMa layer processing. Experimental results show that our proposed CaMa layer helps extract discriminative features that reveal changing regions.
@@ IV Conclusion
Different from previous methods that use CNN and Transformer and mainly focus on spatial change perception modeling, we propose RCaMa to achieve efficient joint spatial-temporal modeling through iterative bi-temporal feature refinement of multiple CaMa layers. Specifically, we introduce Mamba (a state space model) into the RSICC task and propose SD-SSM and TT-SSM. SD-SSM uses differential features to enhance spatial change perception, while TT-SSM promotes bitemporal interactions in a token-wise cross-scanning manner. Experiments demonstrate the superiority of RSCaMa over existing methods, highlighting Mamba’s potential in RSICC. Besides, our systematic evaluation of three language decoding schemes provides valuable insights for future research.